# Practical insights on the effect of different encodings, ansätze and measurements in quantum and hybrid convolutional neural networks

### v1

@@ I.1 Hybrid models I: HQNN-Parallel
This architecture scales really well with the available quantum resources, given that you can tune the number of qubits allocated per circuit. The higher the per circuit qubit count, more correlations between features will be accessible and better performance is expected. Also, given that all circuits are independent, they can be executed in parallel.
In all our experiments for this architecture the feature vectors where of size $768$ and we allocated $8$ qubits per circuit. Both for the $32\times 32\times 3$ and $16\times 16\times 3$ experiments. Therefore, we used a quantum linear layer consisting of $96$ circuits for all one-to-one encodings cases, and $3$ circuits for the amplitude encoding one.
I.2 Hybrid models II: HQNN-Quanv
[…]
Inside each register we apply two identical blocks of a variational ansatz. These blocks are deep enough to mix local features but shallow enough to keep simulation tractable; their parameters, shared across all patches, are the only quantum learnable weights. Upon exit every qubit is measured once in the $X$, $Y$, or $Z$ basis, yielding $Q$ scalars per patch. Aggregating those scalars across the full grid produces $C\times Q$ feature maps of size $H^{\prime}\!\times\!W^{\prime}$, where $H^{\prime}$ and $W^{\prime}$ are the patch indices. Here $C$ is the number of colour channels that are processed independently ($C=3$ for RGB images). A fixed ReLU is applied and the tensor is passed to two classical $3\times 3$ convolutions, global average pooling and a pair of fully connected layers that emit softmax probabilities for the EuroSAT classes.
The learnable weights split into two groups: a small set $P_{\text{q}}$ of rotation angles inside one patch circuit, and the usual classical weights $P_{\text{c}}$ that follow. Because the same circuit is reused for every patch, $P_{\text{q}}$ stays fixed when we move from 32 × 32 to 16 × 16 images or change the stride, while $P_{\text{c}}$ matches the classical head already described for HQNN-Parallel. In practice $P_{\text{q}}$ is only a few hundred parameters; roughly 10–100 times fewer than the weights in a classical first convolution that produces the same number of channel. Yet, it still injects a strong non-linearity at every pixel location.
(a) HQNN-Parallel model from Senokosov et al. (2024).
(b) HQNN-Quanv model inspired by Simonyan and Zisserman (2014).
(c) QCNN model from Cong et al. (2019).
@@ I.2 Hybrid models II: HQNN-Quanv
Figure 4: Scheme of every model used. The three purely quantum models, c), d) and e), are particular cases of the generic PQC structure in Fig. 2. Quantum circuits were created using draw.io, an open-source library Wilkens (2021).
I.3 Pure quantum models I: QCNN
QCNNs were first introduced in Cong et al. (2019). This model performs both the feature extraction and the classification with concatenated quantum operations, so the full model is represented as a unique quantum circuit (see FIG. 4(c)).
It draws inspiration from classical CNN and proposes doing convolutions via 2-qubit gates acting on adjacent qubits; followed by a pooling operation, which reduces the dimensionality by tracing out a subset of the qubits. It had great success due to its remarkable performance given its trainability Hur et al. (2022); Simen et al. (2024).
In this study, we did not want to consider an initial classical stage of dimensionality reduction, since the goal was to check how far we can go with only quantum components. Given limited computational resources, we did not want to exceed the range of 10 to 15 qubits; thus, we used amplitude embedding in order to be able to fit the images in the available number of qubits.
However, if we store the pixels in the amplitudes of the wavefunction, then the analogy of performing a convolution by 2-local gates needs to be looked at again. It turns out that if you track the way in which pixels are mixed in this scenario, you see that pixels from the same vicinity are indeed being mixed together (Appendix A), just like in a convolution. Moreover, by choosing the order in which the amplitudes are fed into the wavefunction, different kernel shapes can be implemented. In this work, we selected several orderings and checked their impact on the QCNN and a simplified model where the effect of the mixing is isolated: the Shaped Embedding Quantum Neural Network (SEQNN).
I.4 Pure quantum models II: SEQNN
@@
This model is used as a baseline, it consists of just a block with the structure described by the Simplified Two Design ansatz, which we used as a fully connected layer in the QCNN architecture.
@@
I.5 Experiments
For each architecture, we systematically evaluate all possible combinations of encoding, ansatz, and measurement methods under identical learning rates, batch sizes, and dataset parameters. This approach aims to isolate and assess the impact of architectural design choices while maintaining consistent hyperparameters. Although further optimization could enhance results, the primary objective of this study is to analyze the effects of these design decisions rather than pursuing performance maximization. See more details in TABLE 1 for the different encodings, ansätze and measurements used for each model.
The dataset comprises of $500$ samples with an $80/20$ split, leading to $50$ examples per class. The hybrid architectures use both $32\times 32$ and $16\times 16$ pixel resolutions. The QCNN model exclusively utilizes $16\times 16$ images to streamline computational requirements while maintaining essential feature extraction capabilities.
@@ I.5 Experiments
- Angle - No entangled - Pauli X The model initially employs
@@
- Angle - No entangled - Pauli X After the quanvolution
@@
I.6 Metrics and optimization
We train our models using the schedule-free variant of the AdamW optimizer Defazio et al. (2024) coupled with the standard Cross Entropy Loss PyTorch (2024). For evaluation, we report classification accuracy throughout the experiments. Additional metrics (e.g., precision, recall, F1-score) were computed using the scikit-learn library Pedregosa et al. (2011), and detailed definitions can be found in its documentation.
@@ II Results
II.1 HQNN-Parallel
By selecting the best validation accuracy results for each encoding-ansätze combination, we observe that several hybrid configurations match or slightly exceed the classical baseline, as shown in FIG. 9. For instance, with $32\times 32$ images, the IQP encoding combined with the no entangled variational circuit achieves $75.1\%$ accuracy, outperforming the classical model’s $67.1\%$.
In FIG. 10, we depict the learning behavior of the best-performing hybrid network and compare it with the classical model. The validation accuracy (FIG. 10c) shows that the best hybrid mode achieves better results than the classical one. However, the validation loss (FIG. 10d) indicates that both models begin to overfit after a few epochs, as the loss starts to increase. Prolonged training would likely exacerbate this issue.
@@ II.1 HQNN-Parallel
Regarding the encoding method, no single solution is universally optimal. However, as seen in FIG. 5, the $RX$, $RY$ and $QAOA$ $Z$ encodings exhibit lower variance in their performance. In contrast, the $Z$ rotation-based encoding ($RZ$) consistently fails, yielding only $10\%$ accuracy. As mentioned in the description of angle encoding, this is due to the encoding only adding an irrelevant global phase.
@@ II.2 HQNN-Quanv
In FIG. 14 depict the best network’s learning behavior and corresponding losses compared with the classic ones. As shown in the validation loss, there is a point at which both the classical and hybrid models cease to learn, leading to an increase in loss. Furthermore, our analysis indicates that the train performance continues to improve during training. However, allowing additional training epochs may lead to overfitting, as prolonged learning could cause the model to excessively adapt to the training data rather than generalizing effectively.
@@
As we can see in FIG. 6, the amplitude and QAOA encodings fail to capture relevant features. In contrast, Waterfall, Ring and Angle encodings demonstrate greater utility for this task, although they yield slightly inferior performance compared to the classical approach. These findings suggest that, when quantum information is available, the aforementioned encodings should be prioritized for implementing quanvolutions of this type. The IQP encoding occupies an intermediate position, indicating potential scenarios where its application could be advantegeous.
@@
In this case we ignored the $Z$ rotations of angle encoding because of the same explained before in section II.A.1
How does the ansätze influence?
Non entangled circuits are less effective than entangled ones. In the FIG. 15(a), the bulk of accuracies for the no entangled ones obtains a worse performance than for the other 3 that do have at least some entanglement. Through this approach, we observe that image features are more effectively captured by leveraging the phenomenon of entanglement.
@@
Prior to conducting the experiments, we expected substantial differences in results depending on the qks, as the circuits employed qks2 qubits. However, it was not the case, we see similar accuracies for qks $=2,3$, maybe a little better for shape 3 (FIG. 16(a)). The results for a qks of size 5 remain pending due to the high computational cost associated with a 25-qubit simulated circuit.
@@ II.3 QCNN
Although our FC model is supposed to mimic an unstructured fully connected layer, it is composed of adjacent two-qubit gates; hence, it is also affected by the change in ordering.
@@
Within each group of measurements used for the quantum architectures, there is still room for some flexibility. In the PauliZ and histogram type of measurements, in some cases there are more active qubits than needed at the end; and in the Paulis, when measuring $N$ qubits, we have in general $14^{N}-$ 1 non-trivial pauli operators to select from. The choice of operator can greatly influence the classification accuracy as well; however, we are picking one at random in the study due to the vast range of possibilities.
In general, the best performing type of measurement was the PauliZ, which required measuring as many qubits as classes. Pauli’s measurement gives the most flexibility at the cost of a great dispersion depending on the chosen operators (FIG. 19).
How does the FC layer deep influence?
We also tested whether increasing the number of parameters of the circuits by making the FC layers deeper influenced the performance of the models. Interestingly, we saw no improvement when increasing the number of parameters for 4 classes, suggesting that architectural changes are needed to enhance model classification capabilities (See FIG. 20). However, for 10 classes, increasing the depth of the FC layer did have a positive impact in the accuracy. A quick look at the training curves (FIG. 18) points towards the fact that we are indeed well in the underparameterization regime (all purely quantum models have less than 400 parameters).
III Evaluation
@@
III.1 Hyperparameters
In this work, we did not consider adjusting the hyperparameters due to the associated computational cost. Future research should explore various combinations of these, such as the number samples used for training, optimizer, learning rate…
Additionally, it would be beneficial to investigate how these hyperparameters affect metrics beyond validation accuracy like generalization and potential barren plateaus. This analysis could provide deeper insights into optimizing our models and improving their performance in practical applications.
III.2 Could we use these models for other types of dataset?
In our study, we utilized the EuroSAT dataset of satellite images Helber et al. (2019). However, this dataset serves more as a means to an end rather than the primary focus of our work. It is suggested that these results may be extrapolated to other types of imagery, potentially yielding slightly better or worse outcomes but exhibiting similar behaviours.
III.3 Position of the quantum layer and effective entanglement

### v2

@@ III.1 Hybrid models I: HQNN-Parallel
This architecture scales really well with the available quantum resources, given that you can tune the number of qubits allocated per circuit. The higher the qubit count per circuit, the more correlations between features will be accessible, and better performance is expected. Moreover, given that all circuits are independent, they can be executed in parallel.
To ensure a consistent dimensionality for the feature across the different input scales, we dynamically adjust the depth of the feature extractor. For $32\times 32\times 3$ inputs, we employ a four-layer convolutional block that reduces the spatial dimensions to $4\times 4$ with $48$ channels. For $16\times 16$ inputs, we use a two-layer block, resulting in an $8\times 8$ feature map with $12$ channels. In both configurations, the resulting flattened feature vector remains constant at $768$ elements. This architectural consistency allows us to utilize a standardized quantum linear layer across all experiments, consisting of $96$ circuits for cases using one-to-one encodings and $3$ circuits for those utilizing amplitude encoding.
III.2 Hybrid models II: HQNN-Quanv
The HQNN-Quanv model follows the high-level blueprint of Senokosov et al. Senokosov2024 but embeds it in the depth hierarchy of VGG16 Simonyan2014, replacing the very first classical convolution with a bank of small, parameter-tied quantum circuits (see FIG. 7(b)). Every input image is split channel-wise into a dense grid of overlapping patches of size $\text{qks}\times\text{qks}$ (we study quantum-kernel size $\text{(qks)}\!=\!2,3$ with stride $1$). Each patch is flattened, normalized, and loaded into a register of $Q=\text{qks}^{2}$ qubits with one of the encodings introduced in Sec. II.2. Because the same circuit parameters are reused for every spatial position and color channel, the layer acts as a genuine quantum analogue of weight-tied convolution.
Inside each register, we apply two identical blocks of a variational ansatz. These blocks are deep enough to mix local features but shallow enough to keep simulation tractable; their parameters, shared across all patches, are the only quantum learnable weights. Upon exit every qubit is measured once in the $X$, $Y$, or $Z$ basis, yielding $Q$ scalars per patch. Aggregating those scalars across the full grid produces $C\times Q$ feature maps of size $H^{\prime}\!\times\!W^{\prime}$, where $H^{\prime}$ and $W^{\prime}$ are the patch indices. Here $C$ is the number of color channels that are processed independently ($C=3$ for RGB images). A fixed ReLU is applied, and the tensor is passed to two classical $3\times 3$ convolutions, global average pooling, and a pair of fully connected layers that emit softmax probabilities for the EuroSAT classes.
The learnable weights split into two groups: a small set $P_{\text{q}}$ of rotation angles inside one patch circuit, and the usual classical weights $P_{\text{c}}$ that follow. Because the same circuit is reused for every patch, $P_{\text{q}}$ stays fixed when we move from $32\times 32$ to $16\times 16$ images or change the stride, while $P_{\text{c}}$ matches the classical head already described for HQNN-Parallel. In practice $P_{\text{q}}$ is only a few hundred parameters, roughly 10–100 times fewer than the weights in a classical first convolution that produces the same number of channels. Yet, it still injects a strong non-linearity at every pixel location.
(a) HQNN-Parallel model from Senokosov2024.
(b) HQNN-Quanv model inspired by Simonyan2014.
(c) QCNN model from Cong2019.
@@ III.2 Hybrid models II: HQNN-Quanv
Figure 7: Scheme of every model used. The three purely quantum models, c), d) and e), are particular cases of the generic PQC structure in Fig. II.2. Quantum circuits were created using draw.io, an open-source library drawio.
III.3 Pure quantum models I: QCNN
QCNNs were first introduced in Cong2019. This model performs both the feature extraction and the classification with concatenated quantum operations, so the full model is represented as a unique quantum circuit (see FIG. 7(c)).
It draws inspiration from classical CNN and proposes doing convolutions via 2-qubit gates acting on adjacent qubits; followed by a pooling operation, which reduces the dimensionality by tracing out a subset of the qubits. It had great success due to its remarkable performance given its trainability Hur2022; simen2024digital.
In this study, we did not want to consider an initial classical stage of dimensionality reduction, since the goal was to check how far we could go with only quantum components. Given limited computational resources, we did not want to exceed the range of 10 to 15 qubits; thus, we used amplitude embedding to be able to fit the images in the available number of qubits.
However, if we store the pixels in the amplitudes of the wavefunction, then the analogy of performing a convolution by 2-local gates needs to be looked at again. It turns out that if you track the way in which pixels are mixed in this scenario, you see that pixels from the same vicinity are indeed being mixed (Appendix A), just like in a convolution. Moreover, by choosing the order in which the amplitudes are fed into the wavefunction, different kernel shapes can be implemented. In this work, we selected several orderings and checked their impact on the QCNN and a simplified model where the effect of the mixing is isolated: the Shaped Embedding Quantum Neural Network (SEQNN).
III.4 Pure quantum models II: SEQNN
@@
This model is used as a baseline; it consists of just a block with the structure described by the Simplified Two Design ansatz, which we used as a fully connected layer in the QCNN architecture.
@@
III.5 Classical Baselines
To strictly isolate the performance impact of the PQCs, we benchmarked each hybrid architecture against a “structural twin” classical baseline. These baselines replace the quantum components with standard classical layers of equivalent dimensionality, while keeping the rest of the architecture identical.
  • In our baseline for HQNN-Parallel, the quantum linear layer branch is bypassed. The feature vector is processed solely by a classical Multi-Layer Perceptron (MLP) consisting of three fully connected layers with decreasing hidden sizes. LeakyReLU activations are applied after the first two layers to introduce non-linearity comparable to the quantum variance.
  • In our baseline for HQNN-Quanv, the initial quantum convolution layer is replaced by a standard classical Conv2d layer. To ensure fair comparison regarding spatial dimension reduction, we employed a kernel size of $3\times 3$, a stride of $2$, and padding of $1$. This configuration matches the effective receptive field and output dimensionality of the quantum kernel approach.
It is important to note that these baselines are controlled architectural references, not attempts to achieve State-of-the-Art (SOTA) results on EuroSAT, which typically requires large-scale pre-trained models.
To ensure the comparison relies solely on architecture rather than optimization tuning, both quantum and classical variants were trained using a centralized trainer with identical hyperparameters. We utilized the AdamWScheduleFree optimizer defazioRoadLessScheduled2024 with a learning rate of $0.01$ and trained for $10$ epochs. Batch sizes were consistent within each experiment type ($16$ for HQNN-Parallel and $4$ for HQNN-Quanv).
III.6 Experiments
For each architecture, we systematically evaluate all possible combinations of encoding, ansatz, and measurement methods under identical learning rates, batch sizes, and dataset parameters. These are benchmarked against the classical baselines defined in the previous section.
This approach aims to isolate and assess the impact of architectural design choices while maintaining consistent hyperparameters. Also, it aims to ensure that performance differences are attributable to the quantum architectural components rather than training variance. Although further optimization could enhance results, the primary objective of this study is to analyze the effects of these design decisions rather than pursuing performance maximization. See more details in TABLE 1 for the different encodings, ansätze and measurements used for each model.
The dataset comprises $500$ samples with an $80/20$ split, leading to $50$ examples per class. The hybrid architectures use both $32\times 32$ and $16\times 16$ pixel resolutions. The QCNN model exclusively utilizes $16\times 16$ images to streamline computational requirements while maintaining essential feature extraction capabilities.
@@ III.6 Experiments
- Angle - No entanglement - Pauli X The model initially employs
@@
- Angle - No entanglement - Pauli X After the quanvolution
@@
III.7 Metrics and optimization
We train our models using the schedule-free variant of the AdamW optimizer defazioRoadLessScheduled2024 coupled with the standard Cross Entropy Loss CEL. For evaluation, we report classification accuracy throughout the experiments. Additional metrics (e.g., precision, recall, F1-score) were computed using the scikit-learn library sklearn, and detailed definitions can be found in its documentation.
@@ IV Results
IV.1 HQNN-Parallel
By selecting the best validation accuracy results for each encoding-ansätze combination, we observe that several hybrid configurations match or slightly exceed the classical baseline, as shown in FIG. 12. For instance, with $32\times 32$ images, the IQP encoding combined with the no entanglement variational circuit achieves $75.1\%$ accuracy, outperforming the classical model’s $67.1\%$.
In FIG. 13, we depict the learning behavior of the best-performing hybrid network and compare it with the classical model. The validation accuracy (FIG. 13c) shows that the best hybrid model achieves better results than the classical one. However, the validation loss (FIG. 13d) indicates that both models begin to overfit after a few epochs, as the loss starts to increase. Prolonged training would likely exacerbate this issue.
@@ IV.1 HQNN-Parallel
Regarding the encoding method, no single solution is universally optimal. However, as seen in FIG. 8, the $RX$, RYRY and QAOAQAOA $Z$ encodings exhibit lower variance in their performance. In contrast, the $Z$ rotation-based encoding ($RZ$) consistently fails, yielding only $10\%$ accuracy. As mentioned in the description of angle encoding, this is due to the encoding only adding an irrelevant global phase.
Interestingly, while QAOAQAOA $Z$ is among the best performing strategies, the encodings QAOAQAOA $X$ and QAOAQAOA $Y$ can fail unpredictably. Numerical analyses (see Appendix F ) suggest that QAOAQAOA $X$ and QAOAQAOA $Y$ have higher expressibility than their analogues with a local field in the $Z$ direction. It has been already established in the field that an excess of expressibility can hamper the trainability of a circuit Larocca2025. Hence, depending on where the optimization is initialized, the optimizer may or may not be able to properly adjust the circuit parameters in order to reach a high accuracy, yielding a big variance in performance, as seen in FIG. 8.
The box plots illustrate the distribution of validation accuracy across the different configurations tested for each category. The central horizontal line indicates the median, the box represents the interquartile range between the 25th and 75th percentiles, and the whiskers extend to the minimum and maximum values excluding outliers, which are plotted as individual dots. Each data point within a category corresponds to a unique combination of encoding and ansatz. While each specific configuration was executed once, preliminary sensitivity analyses confirmed low variance across different parameter initializations, indicating that the reported results are representative of the model’s stable performance.
@@ IV.2 HQNN-Quanv
In FIG. 17 depicts the best network’s learning behavior and corresponding losses compared with the classical ones. As shown in the validation loss, there is a point at which both the classical and hybrid models cease to learn, leading to an increase in loss. Furthermore, our analysis indicates that the train performance continues to improve during training. However, allowing additional training epochs may lead to overfitting, as prolonged learning could cause the model to excessively adapt to the training data rather than generalizing effectively.
@@
As we can see in FIG. 9, the amplitude and QAOA encodings fail to capture relevant features. In contrast, Waterfall, Ring and Angle encodings demonstrate greater utility for this task, although they yield slightly inferior performance compared to the classical approach. These findings suggest that, when quantum information is available, the aforementioned encodings should be prioritized for implementing quanvolutions of this type. The IQP encoding occupies an intermediate position, indicating potential scenarios where its application could be advantageous.
@@
In this case we excluded $Z$ rotations of angle encoding because they only add an irrelevant global phase, as mentioned in section II.A.1 .
How does the ansatz influence?
Non entangled circuits are less effective than entangled ones. In FIG. 18(a), the bulk of accuracies for the non-entangled ones obtains a worse performance than for the other 3 that do have at least some entanglement. Through this approach, we observe that image features are more effectively captured by leveraging the phenomenon of entanglement.
@@
Before conducting the experiments, we expected substantial differences in results depending on the qks, as the circuits employed qks2 qubits. However, it was not the case; we see similar accuracies for qks =2,3=2,3, maybe a little better for shape 3 (FIG. 19(a)). The results for a qks of size 5 remain pending due to the high computational cost associated with a 25-qubit simulated circuit.
@@ IV.3 QCNN
Although our FC model is supposed to mimic an unstructured, fully connected layer, it is composed of adjacent two-qubit gates; hence, it is also affected by the change in ordering.
@@
Within each group of measurements used for the quantum architectures, there is still room for some flexibility. In the PauliZ and histogram types of measurements, in some cases there are more active qubits than needed at the end; and in the Paulis, when measuring $N$ qubits, we have in general $4^{N}-1$ non-trivial Pauli operators to select from. The choice of operator can greatly influence the classification accuracy as well; however, we are picking one at random in the study due to the vast range of possibilities.
In general, the best performing type of measurement was the PauliZ, which required measuring as many qubits as classes. Pauli measurement gives the most flexibility at the cost of a great dispersion depending on the chosen operators (FIG. 22).
How does the FC layer depth influence?
We also tested whether increasing the number of parameters of the circuits by making the FC layers deeper influenced the performance of the models. Interestingly, we saw no improvement when increasing the number of parameters for 4 classes, suggesting that architectural changes are needed to enhance model classification capabilities (see FIG. 23). However, for 10 classes, increasing the depth of the FC layer did have a positive impact on the accuracy. A brief look at the training curves (FIG. 21) points towards the fact that we are indeed well in the underparameterization regime (all purely quantum models have less than 400 parameters).
V Evaluation
@@
V.1 Hyperparameters
In this work, we did not consider adjusting the optimization hyperparameters due to the associated computational cost. Optimizer hyperparameters such as the learning rate, momentum, and adaptivity coefficients can critically affect the geometry of the loss landscape. A large learning rate in high curvature regions may cause oscillations or even divergence, while flat regions can lead to slow (or lack of) convergence if explored with a small learning rate. Similarly, other hyperparameters such as batch size and gradient averaging or momentum control the stochasticity in parameter updates. A loss landscape with many local minima may benefit from a higher effective noise level parameter update, while a smoother loss profile is more efficiently optimized under a more stable parameter update strategy. Additionally, it would be beneficial to investigate how these hyperparameters affect metrics beyond validation accuracy, such as generalization and trainability. Optimizer performance and convergence behavior are inherently landscape-dependent, making systematic hyperparameter sensitivity analysis a nontrivial and computationally demanding task that was out of the scope of the present work.
Regarding trainability, while an empirical analysis of gradient variance was not performed due to current constraints in scaling circuits to a higher number of qubits, existing literature provides a strong theoretical foundation for our approach. Specifically, fundamental studies PhysRevX.11.041011 have analytically demonstrated that QCNN-type architectures are exempt from the Barren Plateau (BP) phenomenon, exhibiting only polynomial gradient vanishing, as long as they maintain logarithmic depth and employ local measurements. For our remaining architectures, we mitigate BP risks by maintaining shallow circuit depths and utilizing local observables, which are recognized strategies to preserve gradient signal in high-dimensional spaces Larocca2025.
V.2 Applicability to Other Datasets
We employed the EuroSAT dataset Helber2019 as a rigorous testbed for our proposed models. EuroSAT was specifically chosen because it offers a necessary balance for NISQ-era quantum simulation: it presents significant classification challenges (high intra-class variance and complex textures) while maintaining a resolution that allows for manageable downsampling without significant feature loss. This makes it superior to simpler datasets, like MNIST variants, for evaluating model performance. Although we anticipate that our results would extrapolate to other types of imagery with similar behaviors, the high computational cost of simulating quantum circuits limits the feasibility of multi-dataset benchmarking in this study. Consequently, the inclusion of further remote sensing datasets remains a priority for future work.
V.3 Position of the quantum layer and effective entanglement